Calculate the pH for each case in the titration of 50.0 mL of 0.100 M HCIO(aq) with 0.100 M KOH(aq). Use the ionization constWhat is the pH after addition of 50.0 mL KOH? pH = What is the pH after addition of 60.0 mL KOH? pH =
Answer #1

PH for weak acide For weak wear acids [+*]= /kac where ka= ionisation constant @ - Concentration. PH for buffer Sowtions : Fo[s*). Skaic - 13x16 € x 1 No3 x10% = 0.547x10 PH = -log[**] = -log (0.547x104) --[ Log (10%)+log (0.547)] --[(-4701 -0.262)No.of mores - Molotily a Volonel nd) 1000 No of initid moles of HUO = 01x50 1000 0.005 OH = 0.12 No.of molek af NOOH = 0.1 xVolume 3. afto addition of 20 m job molek 08 KOH = 0.1x30 = 0.003 1000 - tot Helo 0.005 -0003 F 0.002 toh 0.003 -0.003 Clo thclo tiho e Helo toh 1 C o-005 - + + Co-o05-) К = Снід Сон:] [clo kb can be calculated som Ka&rw Kх Кь: ku -4 -1 го:33х10 = 23pOH = -log Com] = log(41*10) = 5- Log(4.1) = 5-0.61 = 4.39 , pH = 14-poH = 1474.39 = 9.61 5. after addition of 60 ml oH mole